Shift Summary for Evening, 12/Aug/99

Jim Hirschauer (jhirsch@cuhep.Colorado.EDU)
Thu, 12 Aug 1999 23:33:37 -0600 (MDT)

Shift Summary for evening shift 13/Aug/99

On Shift: Ricardo Barbosa, Bob Tschirhart, Jim Hirschauer
Others Present: Julie, Elizabeth, JasonH, Ashkan

Hours of beam delivered at all intensities: 4.0
Hours of beam delivered at nominal intensity: 4.0
Hours of E832 data (special runs not included): 2.0
Hours of special runs (laser scans, muon runs...): 1.0

Accounting of lost time: 2 hrs - CsI HV database value problems

========================================================================
Runs ended this shift: 14206, 14211, 14212
Run 14206 - 7.3M events - BEAM_832
Run 14211 - KTEV_PED
Run 14212 - LASER_SCAN

Tapes brought to FCC this shift: QKC249-QKC254
QKS249-QKS254
QKK249-QKK254
QKB087,QKB088

========================================================================
Accesses during shift: BobT, Julie - to replace divider for 1137, and
DPMT for 2847.

========================================================================
Overview of shift:

Blair Shift Project - there's no way this shift was real

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
15:30 BobT, Ricardo, and JimH take over with no beam,
and CsI HV's all over the place. Julie
has courageously taken control.

16:10 Beam back, but CsI problems persist. PMT gain
drift plot is screwed up. All HV for CsI are
wrong.

17:12 Elliot and Jianbo stop triggers to swap and test
HCC cab B #2 to crate B slot 2. (there is no
problem, this is just a test of board #2, which
seems fine.)

17:30 CsI HV update: Database for SDAQ corrupted. Files
that were generated on 8/10 & 8/11 were corrupted.
HV values look O, but are incorrect. Possible
pointer problem? Luckily Ashkan & Theo had backed
up the SDAQ database on 8/8. Download this file.
Now everything is OK but 1691. JasonH and the
Argonauts loads that one in on time. Now CsI
looks great but CA, HA, COSM, HG, HC are wrong,
distributions & ratio are wrong. Trying to find
the correct database to download. Meanwhile,
Ashkan is still rifling thru backup files.

BUCKLE UP FOR THE IRONY, LADIES AND GENTLEMAN:

Ed calls (& uses the F word) to say that the gain
shift on CsI channel 2084 is due to 1 bad
datapoint. This channel should NOT have its HV
changed.

Someone innocently asks why we messed w/
the HV in the first place.

17:45 Pipeline error - DSP timeout in crate 06. See
logbook for printed copy of ktstat_06.log.
Pipeline reset w/out DBC.

14206 18:10 Run 14206 started.

18:30 BA HV's are wrong for spills 1-18 (~450V too high)

19:21 spill 54 - eerie calm settles on KTeV hall. ValP
starts yelling, "Mayday, Mayday..." "Hey, watch
out..." and then: "Warning, Warning, Warning..."
BobT springs into action is devastated to find
that CsI channel 1137 has left the building.


20:40 Beam goes away, Run 14206 ended.

BobT and Julie begin the first of several
controled accesses. This channel has had
everything swapped previously. Problem must be
arcing or HV surges in the cable plant. We swap
divider with a new divider which has the fuse jumpered.
The channel will still be protected by the current
limit on the LRS 1440 supply. 1137 seems ok.

CsI 2801 found to have a ped shift. We swap the
DPMT 2801, this fixes problem. After swapping
2801, we find a large gain shift in CsI 2847
(10.5%) Could be due to bumping fiber while
working on 2801. Channel seems ok, so we will
change the laser gain.

21:00 KTEV_PED run
Laser scan run
Rosati's Run - delicious

23:30 Machine still down, MI & TEV accesses in progress


========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

o do a muon run when beam comes back

o keep an eye on E/p for CsI channel 2847